# A fast, large-scale optimal transport algorithm for holographic beam shaping

### v1

@@ 1 Introduction
Laser beam shaping is a field with large and growing importance to a wide array of commercial and scientific applications, such as dipole trapping of neutral atoms [1], quantum computing [2], and VR/AR technology [3, 4]. There are many techniques for shaping laser beams [5], but holographic methods stand out for being essentially lossless while allowing arbitrary beam profiles [6, 7]. However, this comes at the cost of a challenging computational problem, known as phase retrieval: One must determine a phase to apply to a laser beam such that the desired target intensity profile is realized in the Fourier plane.
Figure 1: Simulated solution of an example beam shaping problem. All figures are 1024 by 1024 pixels. (a) Input laser beam amplitude profile. The beam is a sum of two Gaussians, with deliberate asymmetry for illustration purposes. (b) Target output beam profile. (c) Optimal transport solution produced by 200 iterations of the FOT algorithm introduced in this work. Solving required 76 seconds and 24 MiB of memory on a Intel Core i7 CPU, or 8.5 seconds on a Nvidia T4 GPU. (d) Solution produced by 100 MRAF iterations initialized from our FOT solution in (c). Solving required a total time (FOT + MRAF) of 94 seconds on an Intel Core i7 CPU, or 9.2 seconds on a Nvidia T4 GPU.
@@
In this work, we exploit special structure of a reformulation of the underlying OT problem to produce an improved algorithm with $\mathcal{O}(N)$ memory scaling. This resolves the main challenge of our previous algorithm, and has the additional benefit of making it much faster. Our new algorithm can solve megapixel scale holographic beam shaping problems in times of order 1 to 100 seconds on an Intel Core i7 CPU, depending on the complexity of the target profile. Problems of such scale were impractical to solve without downsampling in our previous work. An example of the performance of this new algorithm is shown in Fig. 1. Our algorithm is highly parallelizable, and we have demonstrated speedups of 10x on a Nvidia T4 GPU processor, which may allow for wide-ranging real time applications.
@@ 2 Overview in relation to previous work
We recapitulate here the mathematical problem underlying laser beam shaping, as introduced in [7]. To distinguish our new algorithm from previous work, we will refer to the beam shaping algorithm of [7], which relied on a generic optimal transport library, as “Black Box Optimal Transport", or BBOT, and our new algorithm as “Fast Optimal Transport", or FOT. We use lowercase letters to denote quantities associated with the plane of the input laser beam and uppercase letters for corresponding quantities in the Fourier plane. Given are two arrays $g_{jk}$ and $G_{JK}$ representing the discretized amplitudes of the (measured) input laser beam and (desired) target output laser beam, respectively. The values of these arrays are sampled on grids of points $x_{j},y_{k}$ and $X_{J},Y_{K}$, respectively, where $j,k,J,K\in\{0,\dots,n-1\}$ for some integer $n$. We assume square arrays for simplicity, but this is not an essential limitation. Without loss of generality, we may choose units such that the grid coordinate $x_{j}$ has the numerical value $x_{j}=(j-\lfloor n/2\rfloor)/\sqrt{n}$, and similarly for $y_{k},X_{J},Y_{K}$.
@@
where $\mathcal{F}$ is a discrete Fourier transform. Note that $\phi_{jk}$ is a phase measured in cycles rather than radians. The left hand side of this relation represents the amplitude of a Fraunhofer diffraction pattern resulting from applying a phase $\phi$ on an input laser beam $g$. The approximate equality asserts that this pattern should coincide with our target output beam. More rigorously, we typically want to minimize the distance between the left and right hand side of Eq.˜1 with respect to some preferred metric, such as $L^{1}$ or RMS distance.
@@
In BBOT, the OT problem defined by Eqs.˜2 and 3 is solved using a generic OT package [11] as a “black box". While convenient, such a generic OT solver is not able to leverage the special structure of our OT problem, and thus is severely performance limited for large grid sizes $n$. In particular, storing the arrays $\Gamma_{jkLM}$ and $c_{jkLM}$ requires $n^{4}$ floating point numbers each, which for a 1 megapixel image and 64 bit precision becomes 8 terabytes of memory.
@@ 3.1 Convolutional speedup
where $\lambda*V$ or $V*\lambda$ denotes a linear convolution of $\lambda$ with columns or rows of $V$ respectively (see Supplementary Material). The same approach can clearly be applied to Eqs.˜9 and 10 as well. We refer to the variant of Algorithm˜1 using linear convolutions in place of matrix multiplication as “Convolution Fast Optimal Transport", or cFOT, to disambiguate it from FOT.
@@ 4 Complexity Analysis
For a fixed number of iterations the total time complexity of the new algorithm is $\mathcal{O}(n^{3})=\mathcal{O}(N^{3/2})$ for FOT and $\mathcal{O}(n^{2}\log n)=\mathcal{O}(N\log N)$ for cFOT. A theoretical analysis [14] has provided a bound of $\mathcal{O}(\delta^{-2}\log N)$ iterations for convergence to a tolerance $\delta$, as measured by the total variation of marginal constraint violation in the OT problem (see Supplementary Material). Our analysis implies an overall time complexity of $\mathcal{O}(\delta^{-2}N^{3/2}\log N)$ or $\mathcal{O}(\delta^{-2}N\log^{2}N)$ to fixed convergence for the two variants. This is to be compared to $\mathcal{O}(N^{2})$ per iteration and $\mathcal{O}(\delta^{-2}N^{2}\log N)$ to fixed convergence for BBOT. Memory and time complexity for BBOT, FOT, and cFOT is summarized in Table˜1.
Table 1: Asymptotic computational resource requirements of the OT algorithms considered in this work. Here $N=n^{2}$ is the total number of pixels.
Method Memory Time per iteration
BBOT $O(N^{2})$ $O(N^{2})$
FOT $O(N)$ $O(N^{3/2})$
cFOT $O(N)$ $O(N\log N)$
@@ 5 Performance
Fig.˜2 shows scaling of memory and time per iteration versus the total number of pixels of the input and target images, demonstrating the expected $\mathcal{O}(N^{3/2})$ and $\mathcal{O}(N\log N)$ scaling respectively for the FOT and cFOT, along with the steeper $\mathcal{O}(N^{2})$ scaling of BBOT.
Our algorithm has a single hyperparameter $\varepsilon$, which controls the amount of entropic regularization to the OT loss landscape [13]. As shown in Fig.˜3, larger values can provide stronger and faster convergence but have a larger final error. In Fig.˜3(b), the solid and dotted curves show the $L^{1}$ intensity loss (i.e. the $L^{1}$ norm of the difference $\left|\mathcal{F}\left[g_{jk}e^{2\pi i\phi_{jk}}\right]\right|^{2}_{JK}-G^{2}_{JK}$) as a function of $\varepsilon$ for FOT and cFOT phase solutions, respectively. Setting $\varepsilon$ too low can lead to numerical instability in the Sinkhorn-Knopp iterations Eqs.˜7 and 8, as seen in Fig.˜3(b) for cFOT solutions near $\varepsilon\sim 10^{-3}$. We thus typically set $\varepsilon$ to be as low as possible while avoiding such instability. In our FOT implementation, this threshold is typically around $\varepsilon\sim 0.0002$.
@@
Figure 4: Phase solutions (blue) and corresponding output beams (red) from FOT, FOT polished with MRAF, and MRAF initialized with a flat phase. For each output beam, the inset time is the total computation time for generating the phase solution on an Nvidia T4 GPU. The input beam is that shown in Fig.˜1, resized appropriately. FOT was run for 200 iterations with $\varepsilon=0.0002$ and MRAF was run for 100 iterations. For the $n=1024$ and $n=2048$ phases, only the central $256\times 256$ region is shown. Phase vortices in MRAF solutions can be identified by diffraction-limited dark spots in the output beam or corresponding points with non-zero winding number on the phase solution.
@@ 6 Conclusions
We have introduced a new optimal transport-based algorithm for solving phase retrieval problems for holographic beam shaping which relaxes the memory and time complexity constraints of previous work. As with other OT methods, this algorithm is best used in combination with an iterative Fourier transform algorithm or cost function minimization algorithm for subsequent polishing to high solution accuracy. Our new method allows for the solution of realistic megapixel scale phase retrieval problems in timescales of a few seconds and peak memory allocation of tens of megabytes. This performance is sufficient for real time application in some settings, such as dipole trapping of neutral atoms. Another benefit of our method is that it is self-contained, not relying on any black-box OT solvers, which may facilitate its adoption in settings where existing OT packages are not available.

### v2

@@ 1 Introduction
Laser beam shaping is a field with large and growing importance to a wide array of commercial and scientific applications, such as dipole trapping of neutral atoms [1], quantum computing [2], and VR/AR technology [3, 4, 5]. There are many techniques for shaping laser beams [6], but holographic methods stand out for being essentially lossless while allowing arbitrary beam profiles [7, 8]. However, this comes at the cost of a challenging computational problem, known as phase retrieval: One must determine a phase to apply to a laser beam such that the desired target intensity profile is realized in the Fourier plane.
Figure 1: Simulated solution of an example beam shaping problem. All images are 1024 by 1024 pixels. (a) Input laser beam amplitude profile. The beam is a sum of two Gaussians, with deliberate asymmetry for illustration. (b) Target output beam profile. (c) Optimal transport solution produced by 200 iterations of the FOT algorithm introduced in this work. Solving required 76 seconds and 24 MiB of memory on an Intel Core i7 CPU, or 8.5 seconds on a Nvidia T4 GPU. (d) Solution produced by 100 MRAF iterations initialized from our FOT solution in (c). Solving required a total time (FOT + MRAF) of 94 seconds on a CPU or 9.2 seconds on a GPU.
@@
In this work, we exploit special structure of a reformulation of the underlying OT problem to produce an improved algorithm with $\mathcal{O}(N)$ memory scaling. This resolves the main challenge of our previous algorithm, and has the additional benefit of making it much faster. Our new algorithm can solve megapixel-scale holographic beam shaping problems in times of order 1 to 100 seconds on an Intel Core i7 CPU, depending on the complexity of the target profile. Problems of such scale were impractical to solve without downsampling in our previous work. An example of the performance of this new algorithm is shown in Fig. 1. Our algorithm is highly parallelizable, and we have demonstrated speedups of 10x on a Nvidia T4 GPU processor, which may allow for wide-ranging real time applications.
@@ 2 Overview in relation to previous work
We recapitulate here the mathematical problem underlying laser beam shaping, as introduced in [8]. To distinguish our new algorithm from previous work, we will refer to the beam shaping algorithm of [8], which relied on a generic optimal transport library, as “Black Box Optimal Transport”, or BBOT, and our new algorithm as “Fast Optimal Transport”, or FOT. We use lowercase letters to denote quantities associated with the plane of the input laser beam and uppercase letters for corresponding quantities in the Fourier plane. Given are two arrays $g_{jk}$ and $G_{JK}$ representing the discretized amplitudes of the (measured) input laser beam and (desired) target output laser beam, respectively. The values of these arrays are sampled on grids of points $x_{j},y_{k}$ and $X_{J},Y_{K}$, respectively, where $j,k,J,K\in\{0,\dots,n-1\}$ for some integer $n$. We assume square arrays for simplicity, but this is not an essential limitation. Without loss of generality, we may choose units such that the grid coordinate $x_{j}$ has the numerical value $x_{j}=(j-\lfloor n/2\rfloor)/\sqrt{n}$, and similarly for $y_{k},X_{J},Y_{K}$.
@@
where $\mathcal{F}$ is a discrete Fourier transform. Note that $\phi_{jk}$ is a phase measured in cycles rather than radians. The left-hand side of this relation represents the amplitude of a Fraunhofer diffraction pattern resulting from applying a phase $\phi$ on an input laser beam $g$. The approximate equality asserts that this pattern should coincide with our target output beam. More rigorously, we typically want to minimize the distance between the left- and right-hand side of Eq.˜1 with respect to some preferred metric, such as $L^{1}$ or RMS distance.
@@
In BBOT, the OT problem defined by Eqs.˜2 and 3 is solved using a generic OT package [12] as a “black box”. While convenient, such a generic OT solver is not able to leverage the special structure of our OT problem, and thus is severely performance limited for large grid sizes $n$. In particular, storing the arrays $\Gamma_{jkLM}$ and $c_{jkLM}$ requires $n^{4}$ floating point numbers each, which for a 1 megapixel image and 64 bit precision becomes 8 terabytes of memory.
@@ 3.1 Convolutional speedup
where $\lambda*V$ or $V*\lambda$ denotes a linear convolution of $\lambda$ with columns or rows of $V$ respectively (see Supplementary Material). The same approach can clearly be applied to Eqs.˜9 and 10 as well. We refer to the variant of Algorithm˜1 using linear convolutions in place of matrix multiplication as “Convolution Fast Optimal Transport”, or cFOT, to disambiguate it from FOT.
@@ 4 Complexity Analysis
For a fixed number of iterations, the total time complexity of the new algorithm is $\mathcal{O}(n^{3})=\mathcal{O}(N^{3/2})$ for FOT and $\mathcal{O}(n^{2}\log n)=\mathcal{O}(N\log N)$ for cFOT. A theoretical analysis [15] has provided a bound of $\mathcal{O}(\delta^{-2}\log N)$ iterations for convergence to a tolerance $\delta$, as measured by the total variation of marginal constraint violation in the OT problem (see Supplementary Material). Our analysis implies an overall time complexity of $\mathcal{O}(\delta^{-2}N^{3/2}\log N)$ or $\mathcal{O}(\delta^{-2}N\log^{2}N)$ to fixed convergence for the two variants. This is to be compared to $\mathcal{O}(N^{2})$ per iteration and $\mathcal{O}(\delta^{-2}N^{2}\log N)$ to fixed convergence for BBOT. Memory and time complexity for BBOT, FOT, and cFOT is summarized in Table˜1.
Table 1: Asymptotic computational resource requirements of the OT algorithms considered in this work. Here $N=n^{2}$ is the total number of pixels and $\delta$ is convergence tolerance.
Method Memory Time per iteration Time to convergence
BBOT $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(\delta^{-2}N^{2}\log N)$
FOT $\mathcal{O}(N)$ $\mathcal{O}(N^{3/2})$ $\mathcal{O}(\delta^{-2}N^{3/2}\log N)$
cFOT $\mathcal{O}(N)$ $\mathcal{O}(N\log N)$ $\mathcal{O}(\delta^{-2}N\log^{2}N)$
@@ 5 Performance
Fig.˜2 shows scaling of memory and time per iteration versus the total number of pixels of the input and target images, demonstrating the expected $\mathcal{O}(N^{3/2})$ and $\mathcal{O}(N\log N)$ scaling respectively for the FOT and cFOT algorithms, along with the steeper $\mathcal{O}(N^{2})$ scaling of BBOT.
Our algorithm has a single hyperparameter $\varepsilon$, which controls the amount of entropic regularization to the OT loss landscape [14]. As shown in Fig.˜3, larger values can provide stronger and faster convergence but have a larger final error. Fig.˜3(b) and Fig.˜3(c) show the $L^{1}$ intensity loss (i.e. the $L^{1}$ norm of the difference $\left|\mathcal{F}\left[g_{jk}e^{2\pi i\phi_{jk}}\right]\right|^{2}_{JK}-G^{2}_{JK}$) as a function of $\varepsilon$ for FOT and cFOT phase solutions, respectively. Setting $\varepsilon$ too low can lead to numerical instability in the Sinkhorn-Knopp iterations Eqs.˜7 and 8, as seen in Fig.˜3(c) for cFOT solutions near $\varepsilon\sim 10^{-3}$. We thus typically set $\varepsilon$ to be as low as possible while avoiding such instability. In our FOT implementation, this threshold is typically around $\varepsilon\sim 0.0002$.
@@
Figure 4: Phase solutions (blue) and corresponding output beams (red) from FOT, FOT polished with MRAF, and MRAF initialized with a flat phase. For each output beam, the inset time is the total computation time for generating the phase solution on an Nvidia T4 GPU. The input beam is that shown in Fig.˜1, resized appropriately. FOT was run for 200 iterations with $\varepsilon=0.0002$ and MRAF was run for 100 iterations. For the $n=1024$ and $n=2048$ phases, only the central $256\times 256$ region is shown. For the highest resolution target, the rightmost column shows a close-up view at full resolution of a $128\times 128$ pixel region surrounding the bottom right spot of the array. Phase vortices in MRAF solutions can be identified by diffraction-limited dark spots in the output beam or corresponding points with non-zero winding number on the phase solution.
@@ 6 Conclusions
We have introduced a new optimal transport-based algorithm for solving phase retrieval problems for holographic beam shaping which relaxes the memory and time complexity constraints of previous work. As with other OT methods, this algorithm is best used in combination with an iterative Fourier transform algorithm or cost function minimization algorithm for subsequent polishing to high solution accuracy. Our new method allows for the solution of realistic megapixel-scale phase retrieval problems in timescales of a few seconds and peak memory allocation of tens of megabytes. This performance is sufficient for real time application in some settings, such as dipole trapping of neutral atoms. Another benefit of our method is that it is self-contained, not relying on any black-box OT solvers, which may facilitate its adoption in settings where existing OT packages are not available.